OpenStudy (anonymous):

(x – 3)2 + (y + 5)2 = 64 (x + 3)2 + (y + 5)2 = 64 (x – 3)2 + (y – 5)2 = 64 (x + 3)2 + (y – 5)2 = 64

OpenStudy (anonymous):

are you trying to ind x an y

Directrix (directrix):

These look like options for a question that has not been stated. Please state the question. Thanks. (x – 3)2 + (y + 5)2 = 64 (x + 3)2 + (y + 5)2 = 64 (x – 3)2 + (y – 5)2 = 64 (x + 3)2 + (y – 5)2 = 64